### v1

@@ Abstract
The dataset annotation process was based on the hierarchical emotion theory by Shaver et al. [1] and made use of group annotation and majority voting strategies to facilitate label consistency. The resulting dataset contained aspect-level emotion labels for Anger, Sadness, Happiness, Fear, and a None class. Using the new ABEA training dataset, the state-of-the-art ABSA model GRACE by Luo et al. [2] was fine-tuned for ABEA. The results reflected a performance plateau at an F1-score of 70.1% for ATE and 46.9% for joint ATE and AEC extraction. The limiting factors for model performance were broadly identified as the small training dataset size coupled with the increased task complexity, causing model overfitting and limited abilities to generalize well on new data.
@@ 1 Introduction
One of the largest and most widely researched sub-fields of Natural Language Processing (NLP) is Sentiment Analysis (SA). In its most basic form, SA assigns sentiment polarity labels (positive, neutral, and negative) to systematically describe the opinions expressed in text [3, 4]. Beyond basic polarity classification, SA can encompass several other classification tasks, including sarcasm or subjectivity detection [4]. SA can be categorized into different levels based on the scope of the analysis. Document-level SA assigns a single sentiment to an entire document (e.g., article), while sentence-level SA evaluates individual sentences [5]. Recent advancements in computing resources, model architectures, and training data availability have allowed SA research to progress toward a more fine-grained sub-sentence level, so-called aspect-level analysis [6, 7, 8]. It can generate individual outputs for every word or token, where tokens may even be sub-word units [9]. Aspect-based Sentiment Analysis (ABSA) can involve one or more subtasks, which are either addressed sequentially, i.e., individually and then combined to draw conclusions, or jointly as simultaneous outputs of a model [10].
@@
Emotion Analysis (EA) is a sub-discipline of SA that categorizes texts according to distinct emotion classes instead of sentiment polarities. In contrast to SA, the field of EA has yet to undergo a comparable transition to such detailed, sub-sentence level analyses [11, 12, 13, 14]. Given the lack of Aspect-based Emotion Analysis (ABEA) concepts and taxonomies, we follow Zhang et al. [15] for the definitions of the following two subtasks of ABEA (cf. Figure 1): 1) Aspect Term Extraction (ATE): finding the target word(s) of the sentiment in the text, 2) Aspect Emotion Classification (AEC): assigning an emotion label to the respective aspect. In these definitions, an aspect is understood as the target word(s) of an emotion in a given text.
Figure 1: Two sub-tasks of ABEA (adapted from Zhang et al. [15]).
@@
The value of large-scale social media analyses is widely recognized for its direct link to citizen opinions across spatio-temporal dimensions. In ABSA, several studies have incorporated social media data, in particular Twitter data. However, no equivalent studies have emerged thus far for ABEA, leaving a wide research gap with much potential for new insights, methods, and datasets. This gap is particularly relevant for domains such as disaster management, where social media data is increasingly used for real-time monitoring, situational awareness, and response coordination by gauging public needs and opinions [16, 17, 18].
@@
Even studies reviewing the more prominent field of ABSA have highlighted significant dataset-related challenges, such as domain specificity restraining the generalizability of models [19, 20, 10, 21, 22]. Filling the training data gap for ABEA, particularly for social media data, is a prerequisite for any successful model fine-tuning and application. This paper discusses the logistical challenges of generating such a new ABEA training dataset, which encompass the appropriate selection and integration of emotion theory, accounting for human subjectivity [23], and assessing inter-annotator agreement (IAA) at the sub-sentence level. Our second research aim is to implement a model capable of ABEA through adapting the Gradient-Harmonized and Cascaded Labeling for Aspect-Based Sentiment Analysis (GRACE) model by Luo et al. [2]. This endeavor presents various sub-challenges, including the choice of model hyperparameters and methods for performance validation. Consequently, we aim to address the following research questions:
@@ 2.1 Theories of Emotion
An important basis for EA is conceptualizing how emotions are expressed, categorized, and manifested [24, 11, 25, 26]. However, there is no clear consensus on this subject matter, making direct comparisons between research outputs challenging [24, 27]. Originating from the discipline of psychology, two main types of emotion models are commonly used in EA: Categorical models, such as Ekman [28], classify emotions into distinct groups, whereas dimensional models, such as Plutchik [29], position them on continuous scales, such as happiness-sadness or calmness-excitement. Shaver et al. [1] conceptualize a hierarchical emotion graph that incorporates elements of both categorical and dimensional approaches. They identify a set of basic emotions (Love, Joy, Anger, Sadness, Fear, optionally Surprise), which serve as prototypes from which more complex emotions branch out. The closer any two emotion clusters are positioned along the x-axis, the more they are related. Higher y-axis splits indicate stronger sub-branch relationships. This model acknowledges that emotions are not entirely distinct categories but exist on a continuum, with variations within each basic emotion. The hierarchical emotion model [1] therefore forms the theoretical basis for the ABEA model in this study.
@@ 2.2 Training datasets for ABSA & ABEA
Alongside fine-tuning methods, an equally defining factor for the adaptation of pre-trained models for downstream tasks is the availability of task-specific training data. Current research on ABSA is limited by the number of publicly available training datasets and their domain specificity. They are mostly based on customer reviews, which are sourced from e-commerce platforms such as Amazon or Yelp [30]. The most widely used benchmark datasets “SemEval” [31, 32, 33] consist of approximately 3,000 restaurant and laptop reviews and are labeled with aspect terms and sentiment polarity. Alternative ABSA training datasets include “SentiHood” [22], which contains 5,215 sentences with one or more sentiment target terms related to urban neighborhoods, and “MAMS” [34], where each sentence includes multiple aspects and sentiments.
For social media analyses, well-structured datasets like SemEval are unsuitable since they do not resemble the typical complexity, noise, and unstructured nature of social media texts. Three ABSA training datasets consisting of Twitter posts can be found in published literature [35, 36, 37], which are intentionally biased towards named entities like celebrities, companies, or well-known products. In addition, all predominantly contain neutral sentiment labels. Though further social media ABSA training datasets exist, they either do not extract aspect terms [38], or are structured in complex formats like dialogue [39].
Though various datasets for sentence-level emotions have been created [24, 13], ABEA research suffers from a lack of publicly available training datasets annotated with both aspect terms and their corresponding emotions [11] (cf. Table 7). Only the datasets in Maitama et al. [40], De Geyndt et al. [41], Üveges et al. [42] adhere to the definition of aspect terms as the target of an emotion. Among them, only one dataset is entirely in English, and none are publicly available. Furthermore, none of the datasets share the same emotion classification schema. It should also be noted that all datasets are domain-specific, being related to finance and stock markets [43] or Hungarian speeches [42], making them less applicable and effective for general-purpose fine-tuning of ABEA models [10].
@@ 2.3 Models for ABSA
So far, individual ABSA tasks have been more widely researched and have reported higher performance metrics compared to compound tasks [15]. This performance drop is likely due to the differing nature of the sub-tasks, which are difficult to integrate into a single model [2]. For example, ATE and Aspect Sentiment Classification (ASC) are inherently different NLP tasks, the first being a sequence labeling task and the latter being a classification task. Despite this added complexity, unified models offer several advantages over pipeline methodologies, such as their ability to learn co-dependencies and interactions between the individual tasks [2], while also being able to avoid error propagation [44].
The performance of both single-task and compound-task models is most often assessed against the SemEval datasets [31, 32, 33], with most publications reporting the macro F1-score as a means for comparison. In this, single-task ABSA models generally achieve higher performances than compound-task ABSA models. ATE-only and ASC-only models have reported F1-scores up to 89% and 91%, respectively [45, 46]. In comparison, compound ABSA methodologies lag behind in comparative performance, with pair-extraction models reaching F1-scores in the mid-70% to 80% range [47, 48, 49], the highest triplet-extraction models achieving an F1-score of 72% [50, 51, 52], and quad-extraction of 63% [53, 54, 55].
@@ 2.4 Models for ABEA
Many ABEA publications do not focus on or even perform ATE from the input text. Padme and Kulkarni [25] perform sentence-level EA, first identifying general topics using the topic modeling technique Latent Dirichlet Allocation (LDA) and then creating subsets for each topic through keyword filtering. A similar approach is seen in Suciati and Budi [26], where aspect terms are predefined categories and their extraction is not part of the methodology. Instead, they compare how well their methods classify emotions for sentences belonging to the predefined aspect categories. Likewise, Sirisha and Chandana [56] report the use of a hybrid ABSA-RoBERTa-LSTM model to classify Tweets into sentiment and emotion classes without specifying how aspect terms were identified. While Dehbozorgi and Mohandoss [19] seem to extract aspect terms from input text, their method begins by conducting what appears to be sentence-level emotion classification using the Text2Emotion Python package.
@@ 3 Methodology
The methodology presented in this paper consists of two main steps, namely the creation of an ABEA training dataset and the fine-tuning of the GRACE model [2] for ABEA. Figure 2 provides a schematic, high-level overview of our methodology. The code used in this paper is available on GitHub.
@@ 3.1 Data
Social media data from the microblogging platform Twitter (now X) was used to create a ABEA training dataset. A Java-based crawler was employed to collect data, including the Tweet text, a timestamp, and the geolocation, via the v1.1 recent search and streaming Application Programming Interfaces (APIs) [57, 58]. A Tweet’s geolocation was stored either in the form of a point or, more commonly, as a rectangular polygon encompassing a region tagged by the user [59].
Based on the spatial and temporal attributes of the data, an English training dataset consisting of 75% random and 25% disaster-related social media posts was assembled. For this, Tweets from the United Kingdom (UK) and the United States (US) were combined to capture variations in the English language. Tweets for the annotation were then randomly sampled from two larger datasets detailed in Table 1. We incorporated domain-specific training data, assuming that it would contain a higher amount of emotional Tweets and thus addressing the overall class imbalance. The Californian dataset was binary classified with “related” or “unrelated” labels using an XLM-RoBERTa-base model [60] which was fine-tuned specifically for the classification of disaster-relatedness in social media data [61]. Thus, the training dataset was slightly skewed towards disasters, specifically wildfires and floods, potentially increasing the model sensitivity towards emotion and target detection in natural disaster contexts.
@@ 3.2 Training dataset annotation
Due to the lack of publicly available social media ABEA training datasets, a new dataset was annotated (cf. Table 1). In total, seven human annotators from the fields of data science and geoinformatics were enlisted for this task. Each Tweet was labeled by three annotators using an online tool configured with Doccano [62]. The final number of Tweets annotated by three annotators was 2,766. In these Tweets, aspect terms were identified and marked with one of the four basic emotion categories: Happiness, Anger, Sadness, or Fear. For dataset consistency, a fifth, null category None was also included. The annotation followed an iteratively designed guideline, which set out the tools, rules, and processes for annotations. The guideline was tested on a small pilot dataset of 50 Tweets and revised prior to being used for the annotation.
@@ 3.2.1 Emotion detection
For the annotations, two amendments were made to the hierarchical emotion graph by Shaver et al. [1]: Love and Joy were joined into one emotion Happiness, following the suggestion that even Ekman’s basic emotions can be consolidated due to high similarity between certain pairs (e.g., Anger and Disgust) [63]. Furthermore, Surprise was excluded as a stand-alone emotion class as it is a more transient, secondary emotion that triggers another, more dominant one [64]. The adapted graph was provided to the human annotators to help their decision making process.
The inclusion of a neutral class None is common practice in EA to ensure consistency in the annotations, properly handle texts without apparent emotions [65, 66], and ultimately enhance model robustness. This approach reduces computational complexity in identifying aspect terms and prevents overfitting by providing a label for unemotional texts, mitigating bias toward inapplicable emotion classes.
@@ 3.2.3 Inter-annotator agreement
Given the complexity of human language, a high degree of ambiguity related to the annotation of aspect terms was expected. To address the risk of low IAA upfront, the annotators were invited to participate in in-person annotation sessions, where they annotated independently but advanced through the dataset at the same pace, allowing for discussions and clarifications at every Tweet. This method was inspired by Luo et al. [10] and Maitama et al. [40], who employed similar collaborative annotation methods.
All final annotations were determined by means of majority vote [20, 26]. At the character level, the annotated aspect terms of all three annotators were compared and only sequences marked by at least two annotators were kept. If all three annotators agreed that no aspect terms were present, the Tweet was included in the final dataset without aspect term labels. In cases where all three annotators marked aspect terms but none of the labeled sequences overlapped, the Tweet was excluded from the final dataset. A similar method was applied for the final emotion labels, with the additional step that cases without a majority emotion (i.e., all different emotion labels) were manually reviewed. Using this method, 5,704 characters (11%) from the 51,962 initially annotated aspect term characters were removed from the dataset.

### v2

@@ Abstract
The dataset annotation process was based on the hierarchical emotion theory by Shaver.1987 and made use of group annotation and majority voting strategies to facilitate label consistency. The resulting dataset contained aspect-level emotion labels for Anger, Sadness, Happiness, Fear, and a None class. Using the new ABEA training dataset, the state-of-the-art ABSA model GRACE by Luo.2020 was fine-tuned for ABEA. The results reflected a performance plateau at an F1-score of 70.1% for ATE and 46.9% for joint ATE and AEC extraction. The limiting factors for model performance were broadly identified as the small training dataset size coupled with the increased task complexity, causing model overfitting and limited abilities to generalize well on new data.
@@
This is a preprint. For the published, peer-reviewed paper see:
Social Network Analysis and Mining (2026)
https://doi.org/10.1007/s13278-026-01585-5
@@ 1 Introduction
One of the largest and most widely researched sub-fields of Natural Language Processing (NLP) is Sentiment Analysis (SA). In its most basic form, SA assigns sentiment polarity labels (positive, neutral, and negative) to systematically describe the opinions expressed in text [Medhat.2014, Birjali.2021]. Beyond basic polarity classification, SA can encompass several other classification tasks, including sarcasm or subjectivity detection [Birjali.2021]. SA can be categorized into different levels based on the scope of the analysis. Document-level SA assigns a single sentiment to an entire document (e.g., article), while sentence-level SA evaluates individual sentences [Wankhade.2022]. Recent advancements in computing resources, model architectures, and training data availability have allowed SA research to progress toward a more fine-grained sub-sentence level, so-called aspect-level analysis [Torfi.2021, Patwardhan.2023, Liu.2020a]. It can generate individual outputs for every word or token, where tokens may even be sub-word units [Mielke.2021]. Aspect-based Sentiment Analysis (ABSA) can involve one or more subtasks, which are either addressed sequentially, i.e., individually and then combined to draw conclusions, or jointly as simultaneous outputs of a model [Luo.2022].
@@
Emotion Analysis (EA) is a sub-discipline of SA that categorizes texts according to distinct emotion classes instead of sentiment polarities. In contrast to SA, the field of EA has yet to undergo a comparable transition to such detailed, sub-sentence level analyses [DeBruyne.2022, Gupta.2018, Mohammad.2021, MatinPour.2021]. Given the lack of Aspect-based Emotion Analysis (ABEA) concepts and taxonomies, we follow Zhang.2022b for the definitions of the following two subtasks of ABEA (cf. Figure 1): 1) Aspect Term Extraction (ATE): finding the target word(s) of the sentiment in the text, 2) Aspect Emotion Classification (AEC): assigning an emotion label to the respective aspect. In these definitions, an aspect is understood as the target word(s) of an emotion in a given text.
Figure 1: Two sub-tasks of ABEA (adapted from Zhang.2022b).
@@
The value of large-scale social media analyses is widely recognized for its direct link to citizen opinions across spatio-temporal dimensions. In ABSA, several studies have incorporated social media data, in particular Twitter data. However, no equivalent studies have emerged thus far for ABEA, leaving a wide research gap with much potential for new insights, methods, and datasets. This gap is particularly relevant for domains such as disaster management, where social media data is increasingly used for real-time monitoring, situational awareness, and response coordination by gauging public needs and opinions [Yue.2019, Dong.2021, Ma.2014].
@@
Even studies reviewing the more prominent field of ABSA have highlighted significant dataset-related challenges, such as domain specificity restraining the generalizability of models [Dehbozorgi.2021, Orbach.2021, Luo.2022, Tas.2023, Saeidi.2016]. Filling the training data gap for ABEA, particularly for social media data, is a prerequisite for any successful model fine-tuning and application. This paper discusses the logistical challenges of generating such a new ABEA training dataset, which encompass the appropriate selection and integration of emotion theory, accounting for human subjectivity [Zad.2021], and assessing inter-annotator agreement (IAA) at the sub-sentence level. Our second research aim is to implement a model capable of ABEA through adapting the Gradient-Harmonized and Cascaded Labeling for Aspect-Based Sentiment Analysis (GRACE) model by Luo.2020. This endeavor presents various sub-challenges, including the choice of model hyperparameters and methods for performance validation. Consequently, we aim to address the following research questions:
@@ 2.1 Theories of Emotion
An important basis for EA is conceptualizing how emotions are expressed, categorized, and manifested [Acheampong.2020, DeBruyne.2022, Padme.2018, Suciati.2020]. However, there is no clear consensus on this subject matter, making direct comparisons between research outputs challenging [Acheampong.2020, Wang.2020a]. Originating from the discipline of psychology, two main types of emotion models are commonly used in EA: Categorical models, such as Ekman.1992, classify emotions into distinct groups, whereas dimensional models, such as Plutchik.2001, position them on continuous scales, such as happiness-sadness or calmness-excitement. Shaver.1987 conceptualize a hierarchical emotion graph that incorporates elements of both categorical and dimensional approaches. They identify a set of basic emotions (Love, Joy, Anger, Sadness, Fear, optionally Surprise), which serve as prototypes from which more complex emotions branch out. The closer any two emotion clusters are positioned along the x-axis, the more they are related. Higher y-axis splits indicate stronger sub-branch relationships. This model acknowledges that emotions are not entirely distinct categories but exist on a continuum, with variations within each basic emotion. The hierarchical emotion model [Shaver.1987] therefore forms the theoretical basis for the ABEA model in this study.
@@ 2.2 Training datasets for ABSA & ABEA
Alongside fine-tuning methods, an equally defining factor for the adaptation of pre-trained models for downstream tasks is the availability of task-specific training data. Current research on ABSA is limited by the number of publicly available training datasets and their domain specificity. They are mostly based on customer reviews, which are sourced from e-commerce platforms such as Amazon or Yelp [Do.2019]. The most widely used benchmark datasets “SemEval” [Pontiki.2014, Pontiki.2015, Pontiki.2016a] consist of approximately 3,000 restaurant and laptop reviews and are labeled with aspect terms and sentiment polarity. Alternative ABSA training datasets include “SentiHood” [Saeidi.2016], which contains 5,215 sentences with one or more sentiment target terms related to urban neighborhoods, and “MAMS” [Jiang.2019], where each sentence includes multiple aspects and sentiments.
For social media analyses, well-structured datasets like SemEval are unsuitable since they do not resemble the typical complexity, noise, and unstructured nature of social media texts. Three ABSA training datasets consisting of Twitter posts can be found in published literature [Mitchell.2013, Dong.2014, Jiang.2011], which are intentionally biased towards named entities like celebrities, companies, or well-known products. In addition, all predominantly contain neutral sentiment labels. Though further social media ABSA training datasets exist, they either do not extract aspect terms [Crowdflower.2015], or are structured in complex formats like dialogue [Li.2023c].
Though various datasets for sentence-level emotions have been created [Acheampong.2020, Mohammad.2021], ABEA research suffers from a lack of publicly available training datasets annotated with both aspect terms and their corresponding emotions [DeBruyne.2022] (cf. Table 7). Only the datasets in Maitama.2021, DeGeyndt.2022, Uveges.2022 adhere to the definition of aspect terms as the target of an emotion. Among them, only one dataset is entirely in English, and none are publicly available. Furthermore, none of the datasets share the same emotion classification schema. It should also be noted that all datasets are domain-specific, being related to finance and stock markets [Garcia-Mendez.2023] or Hungarian speeches [Uveges.2022], making them less applicable and effective for general-purpose fine-tuning of ABEA models [Luo.2022].
@@ 2.3 Models for ABSA
So far, individual ABSA tasks have been more widely researched and have reported higher performance metrics compared to compound tasks [Zhang.2022b]. This performance drop is likely due to the differing nature of the sub-tasks, which are difficult to integrate into a single model [Luo.2020]. For example, ATE and Aspect Sentiment Classification (ASC) are inherently different NLP tasks, the first being a sequence labeling task and the latter being a classification task. Despite this added complexity, unified models offer several advantages over pipeline methodologies, such as their ability to learn co-dependencies and interactions between the individual tasks [Luo.2020], while also being able to avoid error propagation [Bie.2021].
The performance of both single-task and compound-task models is most often assessed against the SemEval datasets [Pontiki.2014, Pontiki.2015, Pontiki.2016a], with most publications reporting the macro F1-score as a means for comparison. In this, single-task ABSA models generally achieve higher performances than compound-task ABSA models. ATE-only and ASC-only models have reported F1-scores up to 89% and 91%, respectively [Wang.2021d, Wang.2021e]. In comparison, compound ABSA methodologies lag behind in comparative performance, with pair-extraction models reaching F1-scores in the mid-70% to 80% range [Chen.2020, Liang.2021, Cai.2020], the highest triplet-extraction models achieving an F1-score of 72% [Chen.2021a, Xu.2021, Zhang.2021a], and quad-extraction of 63% [Bao.2022, Mao.2022, Zhang.2021b].
@@ 2.4 Models for ABEA
Many ABEA publications do not focus on or even perform ATE from the input text. Padme.2018 perform sentence-level EA, first identifying general topics using the topic modeling technique Latent Dirichlet Allocation (LDA) and then creating subsets for each topic through keyword filtering. A similar approach is seen in Suciati.2020, where aspect terms are predefined categories and their extraction is not part of the methodology. Instead, they compare how well their methods classify emotions for sentences belonging to the predefined aspect categories. Likewise, Sirisha.2022 report the use of a hybrid ABSA-RoBERTa-LSTM model to classify Tweets into sentiment and emotion classes without specifying how aspect terms were identified. While Dehbozorgi.2021 seem to extract aspect terms from input text, their method begins by conducting what appears to be sentence-level emotion classification using the Text2Emotion Python package.
@@ 3 Methodology
The methodology presented in this paper consists of two main steps, namely the creation of an ABEA training dataset and the fine-tuning of the GRACE model [Luo.2020] for ABEA. Figure 2 provides a schematic, high-level overview of our methodology. The code used in this paper is available on GitHub.
@@ 3.1 Data
Social media data from the microblogging platform Twitter (now X) was used to create a ABEA training dataset. A Java-based crawler was employed to collect data, including the Tweet text, a timestamp, and the geolocation, via the v1.1 recent search and streaming Application Programming Interfaces (APIs) [Havas.2021, Schmidt.2023]. A Tweet’s geolocation was stored either in the form of a point or, more commonly, as a rectangular polygon encompassing a region tagged by the user [Honzak.2024].
Based on the spatial and temporal attributes of the data, an English training dataset consisting of 75% random and 25% disaster-related social media posts was assembled. For this, Tweets from the United Kingdom (UK) and the United States (US) were combined to capture variations in the English language. Tweets for the annotation were then randomly sampled from two larger datasets detailed in Table 1. We incorporated domain-specific training data, assuming that it would contain a higher amount of emotional Tweets and thus addressing the overall class imbalance. The Californian dataset was binary classified with “related” or “unrelated” labels using an XLM-RoBERTa-base model [Barbieri.2022] which was fine-tuned specifically for the classification of disaster-relatedness in social media data [Hanny.2024]. Thus, the training dataset was slightly skewed towards disasters, specifically wildfires and floods, potentially increasing the model sensitivity towards emotion and target detection in natural disaster contexts.
@@ 3.2 Training dataset annotation
Due to the lack of publicly available social media ABEA training datasets, a new dataset was annotated (cf. Table 1). In total, seven human annotators from the fields of data science and geoinformatics were enlisted for this task. Each Tweet was labeled by three annotators using an online tool configured with Doccano [Nakayama.2018]. The final number of Tweets annotated by three annotators was 2,766. In these Tweets, aspect terms were identified and marked with one of the four basic emotion categories: Happiness, Anger, Sadness, or Fear. For dataset consistency, a fifth, null category None was also included. The annotation followed an iteratively designed guideline, which set out the tools, rules, and processes for annotations. The guideline was tested on a small pilot dataset of 50 Tweets and revised prior to being used for the annotation.
@@ 3.2.1 Emotion detection
For the annotations, two amendments were made to the hierarchical emotion graph by Shaver.1987: Love and Joy were joined into one emotion Happiness, following the suggestion that even Ekman’s basic emotions can be consolidated due to high similarity between certain pairs (e.g., Anger and Disgust) [Resch.2016]. Furthermore, Surprise was excluded as a stand-alone emotion class as it is a more transient, secondary emotion that triggers another, more dominant one [Ludden.2006]. The adapted graph was provided to the human annotators to help their decision making process.
The inclusion of a neutral class None is common practice in EA to ensure consistency in the annotations, properly handle texts without apparent emotions [DeGeyndt.2023, Plaza-del-Arco.2024], and ultimately enhance model robustness. This approach reduces computational complexity in identifying aspect terms and prevents overfitting by providing a label for unemotional texts, mitigating bias toward inapplicable emotion classes.
@@ 3.2.3 Inter-annotator agreement
Given the complexity of human language, a high degree of ambiguity related to the annotation of aspect terms was expected. To address the risk of low IAA upfront, the annotators were invited to participate in in-person annotation sessions, where they annotated independently but advanced through the dataset at the same pace, allowing for discussions and clarifications at every Tweet. This method was inspired by Luo.2022 and Maitama.2021, who employed similar collaborative annotation methods.
All final annotations were determined by means of majority vote [Orbach.2021, Suciati.2020]. At the character level, the annotated aspect terms of all three annotators were compared and only sequences marked by at least two annotators were kept. If all three annotators agreed that no aspect terms were present, the Tweet was included in the final dataset without aspect term labels. In cases where all three annotators marked aspect terms but none of the labeled sequences overlapped, the Tweet was excluded from the final dataset. A similar method was applied for the final emotion labels, with the additional step that cases without a majority emotion (i.e., all different emotion labels) were manually reviewed. Using this method, 5,704 characters (11%) from the 51,962 initially annotated aspect term characters were removed from the dataset.